Beam commissioning from the Ion Source to Drift Tube Linac 4 was completed this summer, with the following parameters: 62.5 mA, 50 us, 0.2 Hz @74 MeV. Read more here.
After a four-week cooldown period in the Normal Conducting Linac, preparations are now under way to install the fifth and last Drift Tube Linac (DTL5).
Meanwhile in the Superconducting Linac, cryomodule series installation is in progress. With two cryomodules – one medium beta and one spoke model - subject to pilot installation in the spring, placement of the remaining modules needed for Beam on Target, is now in full swing. This was made possible by successful cooldown of the entire cryogenic distribution system, cooldown of the two pilot cryomodules to 2 Kelvin (-271 degrees Celsius) with excellent leak tightness results during vacuum tests.
So far, eight out of the total 13 double-cavity spoke cryomodules for the first section of the Superconducting part, are in place – with connections to waveguides, electricity, water cooling and more ongoing. The modules are welded to the valveboxes of the cryogenic distribution system, with leak tests following.
In the Klystron Gallery, power tests continue for the radio-frequency (RF) power stations and modulators installed in the cells connecting to the linac. All power sources for the spoke and medium beta linac are in place, with the high beta cells currently being populated and energisation following.
Pre-testing of cryomodules continue in the on-site test stand. This involves sending RF power to the cavities, cooled to 2K.
In the last part of the linac – in the Accelerator to Target (A2T) building - installation of quadrupole magnets and the raster system has been performed. Supports are also being placed along the dumpline to the tuning beam dump, with installation of more magnets and a vacuum vessel next up.
All shielding and cooling blocks are in place in the Target Monolith, which is now ready for the installation of major technical components (target wheel, moderator, monitoring plug, instrumentation plug). Piping, connecting from the cooling blocks to the connection ring, is now under way. Connection of these pipes to the cooling systems in the utility building is next, as well as cryogenic piping from the hydrogen room and connected cryostat.
The 42 neutron beam port inserts and plugs are in place around the monolith. These are large steel chunks, containing the first part of the neutron guide for the 15 active instrument ports and the test beamline. Installation of the light shutter system follows. Once installed, all these installations are subject to extensive vacuum testing.
In the target mock-up and test stand, preparations are under way to start the final 24-7 tests of the Target wheel, drive unit and shaft, as well as the Moderator-Reflector Plug and Target Monitoring Plug. The purpose of these tests is to make these technical components ready for installation into the monolith.
The detector system components have been delivered for the LOKI instrument. Read more here.
Infrastructure installations, in and around the cave as well as in-bunker, continue.
Mirror guides have been successfully installed for the ESTIA Selene guide. The in-bunker chopper system has been put in place. The ESTIA cave has been equipped with a local crane and the walls have been painted.
The ODIN cave has been painted white, internally and externally.
In-cave neutron guide has been installed, which means the neutron guide installation for DREAM is substantially completed from bunker to sample position. Infrastructure installations continue inside and outside of the cave.
The BIFROST analyser system is being assembled. The system consists of 45 panels of graphite crystals held with aluminium clips on silicon wafers, mounted on aluminium holders.
Analysers will be mounted on a plate to be installed inside the spectrometer tank. The purpose of the analysers is to reflect the desired energy of neutrons, post-sample, onto the detectors, which are to be mounted at the bottom of the tank.
The analyser system is brought to ESS as a combined In-Kind effort from PSI, DTU and Copenhagen University.
At the same time, at LLB in France, the BIFROST detectors are under extensive testing, with delivery of the installation-ready detector system to ESS expected later this autumn.